OpenStudy (anonymous):

Suppose that △ABC is a right triangle with ∠C=90∘. If AC=4 and BC=6, find the following quantities. cosA sinA tanA secB cscB cotB

OpenStudy (campbell_st):

draw a diagram  when you know AB you can find your trig ratios

OpenStudy (anonymous):

4sqrt3

OpenStudy (anonymous):

cosA=4/4sqrt3?

OpenStudy (campbell_st):

so the hypotenuse AB seems to be incorrect \[AB^2 = 36 + 16...so...AB^2 = 52..... or.... AB = 2\sqrt{13}\] which affects all your answers.

OpenStudy (anonymous):

so cosA=4/2sqrt13?

OpenStudy (campbell_st):

thats correct

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

how would i get the sec csc and cot

OpenStudy (campbell_st):

for remember sec = 1/cos csc = 1/sin cot = 1/tan so find the ratios sin, cos tan then take the reciprocals.

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

but its secB and not secA

OpenStudy (campbell_st):

sin find cosB then take the reciprocal of that answer...

OpenStudy (anonymous):

thanks